Good day, 5Br−(aq)+BrO−3(aq)+6H+(aq)→3Br2(aq)+3H2O(l) The average rate of consumption of Br− is 1.06×10−4 M/s over the first two minutes. The average rate of formation of Br2 during the same time interval is 6.36·10-5 M/s. What is the average rate of consumption of H+ during the same time interval?Thanks
1
Expert's answer
2020-06-11T10:31:38-0400

Rate is proportional to stoichiometric coefficient, so

V(H2) = (6/5)V(Br-) = (6/3)V(Br2) = 1.272*10-4 M/s